OpenStudy (anonymous):

Find the size of each exterior angle in a regular polygon with 12 sides. Would it be 30 degrees?

OpenStudy (angela210793):

yea

OpenStudy (amistre64):

it depends on how you define exterior angle ...

OpenStudy (angela210793):

1800/12=150 so is 30

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=exterior+angle

OpenStudy (amistre64):

180 - (360/12) 180 - 30 = 150 as an exterior angle if your looking for the outside sweep if your looking for the supplement to the interior; then it 180-150 = 30

OpenStudy (amistre64):

there doesnt seem to be a consensus on what "exterior" angles actually measure tho ... so you have to look at the context of the question :)

OpenStudy (angela210793):

ahhhhaaaaaa...thnx :)

OpenStudy (anonymous):

360/12, whatever u call it..:-)